OpenStudy (anonymous):

MEDAL!!! Look at triangle ABC. What is the length of side AB of the triangle?

OpenStudy (anonymous):

we need more info than this

OpenStudy (anonymous):

@Aareon23

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

same 4^2 + 2^2 = AB^2

OpenStudy (anonymous):

6

OpenStudy (danjs):

16 + 4 = AB^2 AB = square root of 20

OpenStudy (anonymous):

yep